# Towards Multimodal Sentiment Analysis Debiasing via Bias Purification

### v1

@@ 1 Introduction
As an essential task in human intention understanding, Multimodal Sentiment Analysis (MSA) [53, 52] attempts to empower machines with the senses of “hearing” [11] and “seeing” [20] to mimic human perception of emotions from diverse modalities. Following the traditional likelihood rule, most existing studies focus on improving MSA performance by exploiting various strategies, including disentangled representation learning [36, 10, 42, 15], attention-based cross-modal interactions [35, 16, 21, 43, 46, 13], fusion mechanisms [26, 51, 19, 29, 32, 44], and well-designed auxiliary tasks [50, 40, 14]. Despite the impressive improvements achieved by numerous works, they all invariably captured harmful dataset biases [31, 22, 6] and suffered from unintended confounders [34, 28, 41], which are multimodal utterance-level label bias and word-level context bias.
@@
Moreover, previous studies [35, 40, 10] have demonstrated that language modality plays an important role in MSA compared to non-linguistic modalities, i.e., a suitable language model could achieve considerable performance [26]. Nevertheless, linguistic information is not always beneficial due to the inherent context bias [17, 28]. The fatal context bias generally emerges when trained models exhibit strong spurious correlations between specific categories and context words in language modality. In Fig. 1(b), some sentiment-irrelevant words appear with imbalanced frequency in negative and positive samples. Consequently, MSA models tend to predict samples containing those words to an incorrect category based on biased statistical information rather than intrinsic textual semantics [39, 28]. For example, Fig. 2(a) shows the predicted binary classification result from a state-of-the-art (SOTA) model [15] on the MOSI. As the context words “good” and “very” appear more frequently in the positive than in the negative samples in the training set, the model predicts the testing sample as “positive” via an unreliable association. Therefore, to perform more reasonable sentiment inference, we need to suitably purify and eliminate the prejudicial effects caused by these biases in prior observations, as shown in Fig. 2(b).
@@
Intuitively, the counterfactual MSA have two outcomes: (1) the trained model relies purely on the statistical shortcut for prediction under the no-treatment condition of the multimodal input. In this case, Fig. 2(b) shows that the purified label bias results in a higher probability of “positive” than “negative”. (2) The trained model relies only on the spurious correlation for prediction under the intervention of preserving context words solely. The result contains the pure side effect obtained by distilling the context bias. Motivated by the above observations, we propose a Multimodal Counterfactual Inference Sentiment (MCIS) analysis framework to mitigate the deleterious impact of two types of dataset biases. Concretely, we first design a tailored causal graph for MSA to diagnose causalities among variables and identify the dataset biases as unintended confounders. The proposed framework is parameter-free and training-free, meaning that MCIS accommodates already-trained models following biased vanilla training via our generalized causal graph. During the inference phase, MCIS intervenes with confounding multimodal inputs via backdoor adjustment theory [23] to mimic the two counterfactual outcomes described above. By subtracting the counterfactual outcomes of the pure dataset biases, MCIS consistently improves the performance of SOTA models with unbiased predictions.
@@ 2 Related Work
Causal Inference. Causal inference is a tool that seeks actual effects in a specific phenomenon [23]. Currently, the mainstream causal inference studies applied to deep learning consist of two aspects: intervention [38, 3, 34] and counterfactuals [22, 33, 31, 28]. Intervention is an operation that alters original data distribution to discover causal effects [7]. Counterfactuals depict imagined outcomes produced by factual variables under different treatments [24]. Our study focuses on obtaining counterfactual outcomes via intervention. Causal inference can remove confounders in data and learn actual causal effects instead of spurious associations, so it has been widely used in many downstream tasks to improve the models’ performance, including visual question answer [22], natural language understanding [34], and scene graph generation [33]. A recent study [31] focused on designing an additional model to capture the harmful effect of textual semantics. However, they ignored the label bias and failed to disentangle the main content and context at the word level, thus incapable of language bias ascription. Different from the above efforts [28, 31], this is the first work to identify both label bias and context bias in the MSA task from a causal perspective. Our framework effectively eliminates the side effects of dataset biases from multimodal inputs, which makes a step towards unbiased prediction in this field.
@@ 3.5 Bias Elimination Strategy
Thanks to humans’ innate counterfactual intuition [22], we can wisely reveal actual causal effects among variables in biased observations rather than superficial connections. The human inference process for unbiased decisions is essentially achieved by comparing factual and counterfactual outcomes [24]. To block the transfer of biases from the training data to the inference process, we imitate such human intuition to propose an operationally simple yet empirically powerful subtraction operation (i.e., bias elimination strategy). The debiased prediction via the strategy is as follows:
@@ 4.3 Implementation Details
Feature Extraction. Following the original protocols of the models above, the audio and visual features are provided by MOSI and MOSEI. The language embeddings are extracted by the pre-trained BERT [4], whether fine-tuning depends on the vanilla settings of different methods. Moreover, we employ the Python NLTK toolkithttps://www.nltk.org/ to tokenize sentences into word lists and then extract the main content words that may affect the semantics in the transcripts. The average mask ratio of the main content words is 68.96%. For the grid search strategy, the search step and search interval are 0.5 and [-2.0, 2.0] in the coarse search process. In the fine search process, the search step is 0.1, while the search interval depends on the results of the coarse search process.
@@ 4.6 Qualitative Analysis
Case Study of Counterfactual Learning. Fig. 5 shows a counterfactual example from each testing set on MOSI and MOSEI, respectively. Specifically, we provide the sentiment intensity scores of positive/negative evaluation results from vanilla DMD, two types of counterfactual outputs, and the counterfactual predictions. The corresponding label and context word distributions for the display samples intuitively show the presence of the dataset biases. Axiomatically, MCIS corrects the baseline predictions and gives reasonable sentiment polarities. Taking Case 1 (Fig. 5(a)) as an example, the vanilla model obtains a falsely positive polarity, which is misled by the dataset biases. According to the two counterfactual outputs corresponding to the purified biases, the biased baseline results suffer from two deleterious effects: (1) the statistical shortcut caused by the large proportion of “positive” labels; (2) the spurious correlation between the context words (e.g., “also”, “very”) and “positive” category. Thanks to the proposed MCIS, we can empower the model to think twice and make unbiased predictions by comparing factual and counterfactual outcomes.
@@ 5 Conclusion
Limitation. Following existing MSA datasets [52, 53], our framework relies on the setting of complete input modalities. When the MCIS-based models are applied to realistic scenarios with missing modalities, both types of counterfactual operations may not adequately capture pure biases, causing sub-optimal gains.

### v2

@@ 1 Introduction
As an essential task in human intention understanding, Multimodal Sentiment Analysis (MSA) [59, 58, 16, 14] attempts to empower machines with the senses of “hearing” [11] and “seeing” [22] to mimic human perception of emotions from diverse modalities. Following the traditional likelihood rule, most existing studies focus on improving MSA performance by exploiting various strategies, including disentangled representation learning [38, 10, 44, 17], attention-based cross-modal interactions [37, 18, 23, 45, 48, 13, 49], fusion mechanisms [28, 57, 21, 31, 34, 46, 50], and well-designed auxiliary tasks [56, 42, 15]. Despite the impressive improvements achieved by numerous works, they all invariably captured harmful dataset biases [33, 24, 6] and suffered from unintended confounders [36, 30, 43], which are multimodal utterance-level label bias and word-level context bias.
@@
Moreover, previous studies [37, 42, 10] have demonstrated that language modality plays an important role in MSA compared to non-linguistic modalities, i.e., a suitable language model could achieve considerable performance [28]. Nevertheless, linguistic information is not always beneficial due to the inherent context bias [19, 30]. The fatal context bias generally emerges when trained models exhibit strong spurious correlations between specific categories and context words in language modality. In Fig. 1(b), some emotionally ambiguous words appear with imbalanced frequency in negative and positive samples. Consequently, MSA models tend to predict samples containing those words to an incorrect category based on biased statistical information rather than intrinsic textual semantics [41, 30]. For example, Fig. 2(a) shows the predicted binary classification result from a state-of-the-art (SOTA) model [17] on the MOSI. As the context words “good” and “very” appear more frequently in the positive than in the negative samples in the training set, the model predicts the testing sample as “positive” via an unreliable association. Therefore, to perform more reasonable sentiment inference, we need to suitably purify and eliminate the prejudicial effects caused by these biases in prior observations, as shown in Fig. 2(b).
@@
Intuitively, the counterfactual MSA have two outcomes: (1) the trained model relies purely on the statistical shortcut for prediction under the no-treatment condition of the multimodal input. In this case, Fig. 2(b) shows that the purified label bias results in a higher probability of “positive” than “negative”. (2) The trained model relies only on the spurious correlation for prediction under the intervention of preserving context words solely. The result contains the pure side effect obtained by distilling the context bias. Motivated by the above observations, we propose a Multimodal Counterfactual Inference Sentiment (MCIS) analysis framework to mitigate the deleterious impact of two types of dataset biases. Concretely, we first design a tailored causal graph for MSA to diagnose causalities among variables and identify the dataset biases as unintended confounders. The proposed framework is parameter-free and training-free, meaning that MCIS accommodates already-trained models following biased vanilla training via our generalized causal graph. During the inference phase, MCIS intervenes with confounding multimodal inputs via backdoor adjustment theory [25, 30] to mimic the two counterfactual outcomes described above. By subtracting the counterfactual outcomes of the pure dataset biases, MCIS consistently improves the performance of SOTA models with unbiased predictions.
@@ 2 Related Work
Causal Inference. Causal inference is a tool that seeks actual effects in a specific phenomenon [25]. Currently, the mainstream causal inference studies applied to deep learning consist of two aspects: intervention [40, 3, 36, 51] and counterfactuals [24, 35, 33, 30, 52]. Intervention is an operation that alters original data distribution to discover causal effects [7]. Counterfactuals depict imagined outcomes produced by factual variables under different treatments [26]. Our study focuses on obtaining counterfactual outcomes via intervention. Causal inference can remove confounders in data and learn actual causal effects instead of spurious associations, so it has been widely used in many downstream tasks to improve the models’ performance, including visual question answer [24], natural language understanding [36], and scene graph generation [35]. A recent study [33] focused on designing an additional model to capture the harmful effect of textual semantics. However, they ignored the label bias and failed to disentangle the main content and context at the word level, thus incapable of language bias ascription. Different from previous efforts [30, 33], this is the first work to identify both label bias and context bias in the MSA task from a causal perspective. Our framework effectively eliminates the side effects of dataset biases from multimodal inputs, which makes a step towards unbiased prediction in this field.
@@ 3.5 Bias Elimination Strategy
Thanks to humans’ innate counterfactual intuition [24, 30], we can wisely reveal actual causal effects among variables in biased observations rather than superficial connections. The human inference process for unbiased decisions is essentially achieved by comparing factual and counterfactual outcomes [26]. To block the transfer of biases from the training data to the inference process, we imitate such human intuition to introduce an operationally simple yet empirically powerful subtraction operation (i.e., bias elimination strategy). The debiased prediction via the strategy is as follows:
@@ 4.3 Implementation Details
Feature Extraction. Following the original protocols of the models above, the audio and visual features are provided by MOSI and MOSEI. The language embeddings are extracted by the pre-trained BERT [4], whether fine-tuning depends on the vanilla settings of different methods. Moreover, we employ the Python NLTK toolkit to tokenize sentences into word lists and then extract the main content words that may affect the semantics in the transcripts. The average mask ratio of the main content words is 68.96%. For the grid search strategy, the search step and search interval are 0.5 and [-2.0, 2.0] in the coarse search process. In the fine search process, the search step is 0.1, while the search interval depends on the results of the coarse search process.
@@ 4.6 Qualitative Analysis
Case Study of Counterfactual Learning. Fig. 5 shows a counterfactual example from each testing set on MOSI and MOSEI, respectively. Specifically, we provide the sentiment intensity scores of positive/negative evaluation results from vanilla DMD, two types of counterfactual outputs, and the counterfactual predictions. The corresponding label and context word distributions for the display samples intuitively show the presence of the dataset biases. Evidently, MCIS corrects the baseline predictions and gives reasonable sentiment polarities. Taking Case 1 (Fig. 5(a)) as an example, the vanilla model obtains a falsely positive polarity, which is misled by the dataset biases. According to the two counterfactual outputs corresponding to the purified biases, the biased baseline results suffer from two deleterious effects: (1) the statistical shortcut caused by the large proportion of “positive” labels; (2) the spurious correlation between the context words (e.g., “also”, “very”) and “positive” category. Thanks to the proposed MCIS, we can empower the model to think twice and make unbiased predictions by comparing factual and counterfactual outcomes.
@@ 5 Conclusion
Acknowledgement. This work is supported in part by the National Key R&D Program of China under Grant 2021ZD0113503 and in part by the Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0103.